OpenStudy (anonymous):

Find the indicated output for this function. If h(x) = x - 2, find h(21).

jimthompson5910 (jim_thompson5910):

Replace x with 21 and evaluate the right side

OpenStudy (anonymous):

whats the answer

jimthompson5910 (jim_thompson5910):

h(x) = x - 2 h(21) = 21 - 2 What's next?

OpenStudy (anonymous):

What is the domain of this relation? {(0, 5), (2, 3), (1, -4), (-3, 3), (-1, -2)}

jimthompson5910 (jim_thompson5910):

The domain is the set of all x values